OpenStudy (anonymous):

Confused help sorry. which type of reasoning is being used in the statement? I know that 4(3x + 2) = 12x + 8 because of the Distributive Property.Inductive,or Deductive

Directrix (directrix):

I say deductive reasoning because: Distributive Property: For a, b, and c: a(b+c) = ab +ac Therefore, 4(3x + 2) = 12x + 8

OpenStudy (anonymous):

that makes sense thank you again :)

Directrix (directrix):

@ lalaland --> Asking a question on OS-Math means never having to say you are sorry.

OpenStudy (anonymous):

ok thank you I just dont want to bother